OpenStudy (anonymous):

z^2-7z+10 ---------- simplify by removing factors of 1 z^2+6z-16

OpenStudy (anonymous):

\[=(z-5)(z-2)/(z+8)(z-2)\] \[=(z-5)/(z+8)\]